Bill is three times older than jane. If bill were 9 years younger and jane were 6 years older, their ages would differ by 21 years. How old is Jane
Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

Wrte an equation for each statement
:
Bill is three times older than jane.
b = 3j
If bill were 9 years younger and jane were 6 years older, their ages would differ by 21 years.
(b-9) - (j+6) = 21
b - 9 - j - 6 = 21
b - j - 15 = 21
b - j = 21 + 15
b - j = 36
Replace b with 3j (from the 1st statement)
3j - j = 36
2j = 36
j = 18 yrs is Jane's age
:
:
you can find b, and check solutions in the 2nd statement

Answer by Alan3354(69443)

Bill is three times older than jane. If bill were 9 years younger and jane were 6 years older, their ages would differ by 21 years. How old is Jane
-------------
3 times older = 4 times as old.
---
1 time older = 100% older = 2 times as old
2 times older = 200% older = 3 times as old
3 times older = 300% older = 4 times as old
==============================
B = 4J
B-9 = J+6 + 21
4J-9 = J+27
3J = 36
J = 12 years old
B = 48 years old
-------------
Is Bill a Republican in Alabama?